A bottle of juice drink is claimed to be 500 mL with standard deviation of 20 mL. Ramil wants to investigate the same fruit juice drink to estimate the true population mean content correct to within 5mL, adopting a 95% confidence. How many bottles of the fruit juice drink of the same brand does he need for a sample?


1
Expert's answer
2021-05-13T06:20:51-0400

"ME=5ml"

"\:z-score\:corresponding\:to\:95\%\:confidence=1.96\:"

"\sigma =20ml"

"sample\:size,\:n=\left(\:\frac{z\sigma \:}{ME}\right)^2"

"=\left(\:\frac{1.96\times 20\:}{5}\right)^2"

"=61.4656\approx 62\:bottles\:of\:juice"